how do you find the length and the midpoint of a segment using two endpoints:A'
(2,-7) and B'(-5,-4)

Answer by rfer(16322)

I used the Pythagorean theorem
a^2+b^2=c^2
3^2+7^2=c^2
9+49=58
c=7.6
so minpoint is 7.6/2=3.8